A 455-foot fence encloses a pasture. What is the length of each side of the pasture? The shape of the pasture is a trapezoid (sort of).
One side says 3x ft, one side says 1.5x ft, one side says 180ft, and the last side says x ft.
How do I solve this problem?

Answer by ikleyn(52781)

.

The perimeter is the sum of all sides


   455 = 3x + 1.5x + 180 + x


which is the same as


    455 = 5.5x + 180,

    455 - 180 = 5.5x

      275     = 5.5x

        x     = 275/5.5

        x     = 50 ft.


Therefore, one side is 3*50 = 150 ft;  other side is  1.5*50 = 75 ft;  the last side is  50 ft.    ANSWER

Solved and explained.